OpenStudy (anonymous):

Let f(t)=8t2+4t+1 . Find the value of t for which the tangent line to the graph of f(t) has slope 1

OpenStudy (chihiroasleaf):

do you mean \[f(t)=8t^2+4t+1\] ?

OpenStudy (anonymous):

yes

OpenStudy (chihiroasleaf):

do you know how to find the slope of a tangent line of a function in a point?

OpenStudy (anonymous):

do you just use y=mx+b?

OpenStudy (anonymous):

you should take derivative to get f'(t) and then let it =1 to solve for t. Hope this help

OpenStudy (anonymous):

f'(t) = 16t +4 = 1 ---> t = -3/16

OpenStudy (anonymous):

thank you!